You are planning to save for retirement over the next 20 years. To do this, you will invest $1,000 a month in a stock account and $700 a month in a bond account. The return of the stock account is expected to be 11 percent, and the bond account will pay 7 percent. When you retire, you will combine your money into an account with a 10 percent return.


Required:

How much can you withdraw each month from your account assuming a 20-year withdrawal period?(Do not round your intermediate calculations.)
Answer #1

Please provide rating..

we have to use financial calculator to solve this problem
First we have to compute the future value at 20th year.
Stock investment future value
Put in calculator
PV 0
PMT -1000
I =11%/12 0.92%
N=20*12 240
Compute FV $865,638.04
Bond investment future value
Put in calculator
PV 0
PMT -700
I =7%/12 0.58%
N=20*12 240
Compute FV $364,648.66
Total future value at the end of 20th year = $1,230,286.70
We have compute next 20 year annuity withdrawal
Put in calculator below
PV ($1,230,286.70)
I=10%/12 0.83%
N=20*12 240
FV= 0
Compute PMT $11,872.53
hence, we can withdraw $11872.53 per month .